OpenStudy (anonymous):

A motion is described as \(v^2=-4x^2+8x+4 \), started at \(x=0\). find the \(x-t\) equation which is the displacement equation.

OpenStudy (anonymous):

This is what I did: \[v^2=-4x^2+8x+4 => v= \sqrt{-4x^{2}+8x+4} => \frac{dt}{dx} = \frac{1}{\sqrt{-4x^{2}+8x+4}} \] \[=> t = \int\limits\frac{1}{\sqrt{-4x^{2}+8x+4}} dx => t = \frac{1}{2} \sin^{-1} \left(\frac{x-1}{\sqrt{2}}\right) -\frac{\pi}{8} \] then when I solve for \(x\) I will get an ugly result. So, I think that I did it wrong to obtain the \(x-t\) equation..

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

then what is the problem

OpenStudy (anonymous):

correct? lol I think it is wrong; since when i solve for \(x\) it won't look nice. And i have to graph it later.. I think I did it wrong

OpenStudy (amistre64):

im wondering if you have to undo the square to do this, or if you can just integrate both sides and call it a day

OpenStudy (anonymous):

integrate both sides? using the acceleration?

OpenStudy (amistre64):

displacement is the integration of velocity right?

OpenStudy (anonymous):

yeah.. or do i use this..since you said integrate both sides\[\frac{dv}{dt} =\frac{d}{dx} \left(\frac{1}{2}v^{2}\right) \]

OpenStudy (anonymous):

wait displacement is the differentiate of velocity i thinnk

OpenStudy (anonymous):

differentiation*

OpenStudy (amistre64):

\[\int (v^2= -4x^2 +8x +4)\] \[\frac{1}{3}v^3= -\frac{1}{3}4x^3 +\frac{1}{2}8x^2 +4x +C\] is what i was thinking, could be off tho :)

OpenStudy (amistre64):

displacement is integration of velocity; the derivative of velocity is acceleration

OpenStudy (anonymous):

isn't the displacement \(x\) is the differentiation of velocity since \(\frac{dx}{dt}\) is the velocity? And I am not sure about your method..

OpenStudy (amistre64):

im not to sure about it either. I might not be remembering the correct definition of "differentation". I might be confusing it with taking a derivative. At any rate, integrating velocity is akin to determining displacement.

OpenStudy (anonymous):

well, when you integrate the velocity you get the displacement i think that you're right

OpenStudy (amistre64):

v^2 = -4x^2 +8x +4 v^2 = -4(x^2 -2x -1) v^2 = -4(x^2 -2x +1 -1-1 ) v^2 = -4 [(x-1)^2 -2] v^2 = -4(x-1)^2 +8

OpenStudy (anonymous):

yeah i did that to integrate it..

OpenStudy (amistre64):

im curious if the domain is relevant. just thoughts .... v^2 is always positive, and the parabola is hill shaped so that restricts the domain; and also the x >= 0 part hmmm

OpenStudy (anonymous):

i don't think that it's relevant in this question..and \(x=0\) is given anyway then its \[v= \sqrt{-4x^{2}+8x+4} \]

OpenStudy (amistre64):

as long as you did the integrating part correctly :) i cant see anything else to worry about

OpenStudy (anonymous):

I am not sure if i did the correct method; and i am not sure how to solve for \(x\) if i did it correctly

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+sqrt%28-4x%5E2%2B8x%2B4%29 gives the integration of sqrt(-4x^2 +8x +4) and yeah, it does look a bit messy.

OpenStudy (amistre64):

i wish i knew more about wave equations to be able to be useful ....

OpenStudy (anonymous):

ok, thanks anyway!